Rosa is three times as old as her daughter Maria. The sum of their ages is 44 years. How old is Maria?
Answer by mananth(16946)

Maria = x years
Rosa =3x years
...
sum = 44
..
3x+x=44
4x=44
...
divide both sides by 4
4x/4 = 44/4
x=11 Maria's age
Rosa's age = 44-11
33 years
...
m.ananth@hotmail.ca